OpenStudy (thomas5267):

Prove: If p is prime and \(p|c^2\) then \(p|c\) and \(p^2|c^2\).

OpenStudy (thomas5267):

Can I argue that \(c\) factors into unique primes \(p_1^{a_1}p_2^{a_2}\ldots p_n^{a_n}\)and \(c^2=p_1^{2a_1}p_2^{2a_2}\ldots p_n^{2a_n}\) so if \(p|c^2\) then \(p|c\) as the exponent are all even and by extension \(p^2|c^2\)?

OpenStudy (thomas5267):

Please check my prove. @ganeshie8

ganeshie8 (ganeshie8):

that works but i think a proof without calling for prime factorization gives more insight into other properties...

ganeshie8 (ganeshie8):

Save prime factorization proof, I would try proof by contradiction as it tells me what I need to start with atleast..

ganeshie8 (ganeshie8):

Prove If \(p|c^2\) then \(p|c\) and \(p^2|c^2\) Proof (by contradiction) : Suppose \(p|c^2\) but \(p \nmid c\), then we have \(c = pn+r\) for some \(0\lt r \lt p\) Since \(p|c^2\), we have \(c^2 = pm\) for some integer \(m\) \(\implies (pn+r)^2 = pm\) \(\implies p(pn^2 + 2nr)+r^2 = pm\) which is impossible unless \(r=0\). contradiction. \(\blacksquare\)

ganeshie8 (ganeshie8):

Contradiction occurs because we have assumed \(r\) is non zero which is required for \(p \nmid c\)